OpenStudy (anonymous):

Simplify the following expression: (x + 6y) - (3x − 10y). If the final answer is written in the form Ax + By, what is the value of A? Numerical Answers Expected! I think A is -2 am I correct?

OpenStudy (anonymous):

@whpalmer4 am I right?!

OpenStudy (anonymous):

x+6y-3x+10y=-2x+16y, so yes A=-2 and B=?

OpenStudy (whpalmer4):

yes!

OpenStudy (anonymous):

(x + 6y) - (3x − 10y) =x-3x +6y +10y = -2x +16y comparing it with Ax + By we get A = -2 and B= 16

OpenStudy (anonymous):

@dpasingh & @Umangiasd I got -16 for B! & thanks @whpalmer4

OpenStudy (whpalmer4):

oh, not so good.... 6 -(-10) = ?

OpenStudy (whpalmer4):

\[(x+6y)-(3x-10y)\]\[(x+6y)-1(3x-10y) = x+6y -1*3x -1*(-10)y \]\[= x+6y-3x+10y = -2x+16y\]

OpenStudy (anonymous):

its not -2 ? @whpalmer4

OpenStudy (whpalmer4):

no, your value for B is incorrect.

OpenStudy (anonymous):

okay I made a mistake I had -4 as the value of B.. lol @whpalmer4 not -16!

OpenStudy (whpalmer4):

you wrote: "@dpasingh & @Umangiasd I got -16 for B!"

OpenStudy (anonymous):

yeah! I just realized lol! But I worked it out on paper & I got -4! & I posted this after I worked it out.. I was looking at a different problem. thats all. @whpalmer4

OpenStudy (whpalmer4):

okay. as long as you understand that A = -2, B = 16 is the only correct answer for this problem.

OpenStudy (anonymous):

okay! thanks! :) @whpalmer4